# Uniqueness of free 2-periodicities of links

### v1

@@ 4. Prime and JSJ decompositions
In this section, we describe the prime and JSJ decompositions of the complement of a link in $\mathbb{R}\mathbb{P}^{3}$, and we prove the main theorem.
The prime decomposition of a compact orientable 3-manifold is the unique maximal decomposition by the connected sum. Let $L$ be a link in a compact orientable irreducible 3-manifold $X$. The link $L$ is called a split link if the complement of $L$ is reducible. We say that $L$ splits to links $L^{\prime},L_{1},\dots,L_{m}$ if $L=L^{\prime}\sqcup L_{1}\sqcup\dots\sqcup L_{m}$ and there are disjoint balls $B_{1},\dots B_{m}$ in $X$ disjoint from $L^{\prime}$ such that each $B_{i}$ contains $L_{i}$. Then a split summand $L_{i}$ can be regarded as a link in $S^{3}$. If $X\neq S^{3}$ and $m\geq 1$, then $L$ is a split link. The prime decomposition implies the following lemma.
@@
The JSJ decomposition of a compact orientable irreducible 3-manifold with toral boundary is the unique minimal decomposition along disjoint (possibly empty) incompressible tori (called the JSJ tori) such that each piece after decomposition is Seifert fibered or (algebraically) atoroidal. Thurston’s geometrization [27] implies that an atoroidal 3-manifold with toral boundary admits a hyperbolic metric of finite volume. Consequently, the complement of a non-split link in $X$ admits a unique decomposition along tori into pieces each of which is Seifert fibered or hyperbolic.
To consider the JSJ decomposition of the complement of a link in $\mathbb{R}\mathbb{P}^{3}$, we need describe a torus in $\mathbb{R}\mathbb{P}^{3}$. A knotted hole ball in a 3-manifold $X$ is the complement of a properly embedded arc in a 3-ball contained in $X$. Note that the complement of a knotted hole ball in $X$ is the connected sum of $X$ and $S^{1}\times D^{2}$. A knotted hole ball may be a solid torus.
@@
Since $\pi_{1}(\mathbb{R}\mathbb{P}^{3})\cong\mathbb{Z}/2\mathbb{Z}$ is finite, any torus $T$ in $\mathbb{R}\mathbb{P}^{3}$ is compressible. A sphere $S$ is obtained by compressing $T$ along a compressing disk $D$. The sphere $S$ bounds a ball $B$ since $\mathbb{R}\mathbb{P}^{3}$ is irreducible. If $B$ and $D$ are disjoint, then $T$ bounds a solid torus. Otherwise, $T$ bounds a knotted hole ball. ∎
Let $L$ be a non-split link in $\mathbb{R}\mathbb{P}^{3}$. we say that a JSJ piece $M$ of the complement of $L$ is outermost if $\overline{\mathbb{R}\mathbb{P}^{3}\setminus M}$ is a disjoint union of solid tori and knotted hole balls. This definition can be applied to a non-split link in $S^{3}$ or $S^{1}\times D^{2}$. However, every JSJ piece for a non-split link in $S^{3}$ is outermost. The outermost piece for a non-split link in $S^{1}\times D^{2}$ contains the boundary of $S^{1}\times D^{2}$, and so it is unique. In any case, an outermost piece can be reembedded to a link complement by reattaching the 1-handles adjacent to the knotted hole balls as in [9, Proposition 2.2] and [21, Lemma 3.6].
@@
  • If $T$ bounds two solid tori (i.e. $T$ is a Heegaard torus), then $e$ is unoriented.
@@
Since $G_{L}$ is a tree, such a function $f$ exists. Suppose that a vertex $v\in V_{L}$ attains a local minimum of $f$. Then each edge incident to $v$ is oriented from $v$ or unoriented. Let $M$ denote the JSJ piece corresponding to $v$. Each boundary component of $M$ is a JSJ torus or the boundary of neighborhood of a component of $L$. The orientations of edges imply that each boundary component of $M$ bounds a solid torus or a knotted hole ball disjoint from $M$. Hence $M$ is outermost. Since $G_{L}$ is finite, the minimum of $f$ exists. ∎
For a non-split link $L\subset\mathbb{R}\mathbb{P}^{3}$, let $\widetilde{L}\subset S^{3}$ denote the preimage of $L$ by the covering map $\pi\colon S^{3}\to\mathbb{R}\mathbb{P}^{3}$. By [21, Lemma 2.10] and the fact that there exist no embedded Klein bottles in $\mathbb{R}\mathbb{P}^{3}$ [8, 17], the JSJ tori of the complement of $\widetilde{L}$ is the preimage of the JSJ tori of the complement of $L$ by the covering map $\pi$. An outermost JSJ piece for $L$ is characterized as follows.
@@
Note that the preimage of a knotted hole ball in $\mathbb{R}\mathbb{P}^{3}$ consists of its two copies. The preimage of a solid torus in $\mathbb{R}\mathbb{P}^{3}$ is one or two solid tori.
Suppose that $M$ is outermost. In other words, each component of $\overline{\mathbb{R}\mathbb{P}^{3}\setminus M}$ is a solid torus or a knotted hole ball. If $\widetilde{M}$ is disconnected, then the preimage of each component of $\overline{\mathbb{R}\mathbb{P}^{3}\setminus M}$ is disconnected, and their union $S^{3}$ would be also disconnected. Hence $\widetilde{M}$ is connected. Since the components of $\overline{S^{3}\setminus\widetilde{M}}$ that are not solid tori are the preimage of such components of $\overline{\mathbb{R}\mathbb{P}^{3}\setminus M}$, their number is even.
Conversely, suppose that $M$ is not outermost. Then Lemma 4.2 implies that $M$ is contained in a solid torus or a knotted hole ball. If $M$ is contained in a knotted hole ball, the preimage $\widetilde{M}$ is disconnected. Suppose that there is a JSJ torus $T\subset\partial M$ such that $M$ is contained in a solid torus $S$ bounded by $T$, and $T$ is not a Heegaard torus. If $\widetilde{M}$ is connected, then the components of $\overline{S^{3}\setminus\widetilde{M}}$ that are not solid tori consist of
@@
and so their number is odd. ∎
@@
The argument for split links is identical to that in the proof of [21, Theorem 4.5]. Suppose that a link $L_{0}$ maximally splits to $L_{0}^{\prime},L_{0,1},\dots,L_{0,m}$, where $L_{0}^{\prime}$ is a (possibly empty) non-split link. Then $\widetilde{L_{0}}=\pi^{-1}(L_{0})$ maximally splits to $\widetilde{L_{0}^{\prime}}=\pi^{-1}(L_{0}^{\prime})$ and $L_{0,j}^{k}$ for $1\leq j\leq m$ and $k=1,2$, where $L_{0,j}^{k}$ is a copy of $L_{0,j}$ as a link in $S^{3}$. The same applies to $L_{1}$. Since $\widetilde{L_{0}}$ and $\widetilde{L_{1}}$ are isotopic, their split summands are isotopic by Lemma 4.1. Hence we may assume that $L_{0}$ and $L_{1}$ are non-split links.
We consider the JSJ decompositions of the complements of $L_{0}$, $L_{1}$, and $\widetilde{L_{0}}\cong\widetilde{L_{1}}$. As mentioned above, the JSJ decomposition of the complement of $\widetilde{L_{0}}\cong\widetilde{L_{1}}$ is the lifts of those of $L_{0}$ and $L_{1}$. By Lemma 4.3, there is an outermost JSJ piece $M_{0}$ of the complement of $L_{0}$. By Lemma 4.4, $\widetilde{M_{0}}=\pi^{-1}(M_{0})$ is a single JSJ piece of the complement of $\widetilde{L_{0}}$. There is a JSJ piece $M_{1}$ of the complement of $L_{1}$ such that $\widetilde{M_{1}}=\pi^{-1}(M_{1})$ is isotopic to $\widetilde{M_{0}}$. By Lemma 4.4, the piece $M_{1}$ is also outermost. Since $M_{0}$ and $M_{1}$ are outermost, they can be reembedded to the complements of links in $\mathbb{R}\mathbb{P}^{3}$. Since the preimage of a knotted hole ball in $\mathbb{R}\mathbb{P}^{3}$ consists of its two copies in $S^{3}$, the pieces $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$ are equivariantly reembedded to link complements. Since $M_{0}$ and $M_{1}$ are hyperbolic or Seifert fibered, they are isotopic in $\mathbb{R}\mathbb{P}^{3}$ by Theorems 2.5 and 3.6.
We show that the isotopy between $M_{0}$ and $M_{1}$ extends to an isotopy between the complements of $L_{0}$ and $L_{1}$ using an isotopy between the complements of $\widetilde{L_{0}}$ and $\widetilde{L_{1}}$. Since $M_{0}$ and $M_{1}$ are outermost, the remaining regions consist of solid tori and knotted hole balls. Since the preimage of a knotted hole ball in $\mathbb{R}\mathbb{P}^{3}$ consists of its two copies in $S^{3}$, the isotopy extends to the knotted hole balls. By [21, Theorem 4.5] for links in the solid torus, the isotopy in $S^{3}$ induces an isotopy between links in the remaining solid tori contained in $\mathbb{R}\mathbb{P}^{3}$. Therefore $L_{0}$ and $L_{1}$ are isotopic. ∎
@@ Acknowledgements
The author is grateful to Yuya Koda, Yuka Kotorii, Sonia Mahmoudi, Elisabetta Matsumoto, and Yuta Nozaki for their helpful discussions. This work is supported by the World Premier International Research Center Initiative Program, International Institute for Sustainability with Knotted Chiral Meta Matter (WPI-SKCM2), MEXT, Japan.

### v2

@@ 4. Prime and JSJ decompositions
In this section, we describe the prime and JSJ decompositions of the complement of a link in $\mathbb{R}\mathbb{P}^{3}$, and we prove the main theorem. See [19] for details of the prime and JSJ decompositions.
The prime decomposition [23, 27] of a compact orientable 3-manifold is the unique maximal decomposition by the connected sum. Let $L$ be a link in a compact orientable irreducible 3-manifold $X$. The link $L$ is called a split link if the complement of $L$ is reducible. We say that $L$ splits to links $L^{\prime},L_{1},\dots,L_{m}$ if $L=L^{\prime}\sqcup L_{1}\sqcup\dots\sqcup L_{m}$ and there are mutually disjoint balls $B_{1},\dots B_{m}$ in $X$ disjoint from $L^{\prime}$ such that each $B_{i}$ contains $L_{i}$. Then a split summand $L_{i}$ can be regarded as a link in $S^{3}$. If $X\neq S^{3}$ and $m\geq 1$, then $L$ is a split link. The prime decomposition implies the following lemma.
@@
The JSJ decomposition [21, 22] of a compact orientable irreducible 3-manifold with toral boundary is the unique minimal decomposition along mutually disjoint (possibly empty) incompressible tori (called the JSJ tori) such that each piece after decomposition is Seifert fibered or (algebraically) atoroidal. Thurston’s geometrization [34] implies that an irreducible, boundary-irreducible, atoroidal 3-manifold with non-empty toral boundary admits a hyperbolic metric of finite volume. Consequently, the complement of a non-split link in $X$ admits a unique decomposition along tori into pieces each of which is Seifert fibered or hyperbolic.
To consider the JSJ decomposition of the complement of a link in $\mathbb{R}\mathbb{P}^{3}$, we need to describe a torus in $\mathbb{R}\mathbb{P}^{3}$. We call a submanifold $Y$ of a 3-manifold $X$ that satisfies the following conditions a knotted hole ball.
  • There is a closed 3-ball $B$ contained in $X$.
  • There is an embedding of a 1-handle $\iota\colon[0,1]\times D^{2}\to B$.
  • $\iota^{-1}(\partial B)=\{0,1\}\times D^{2}$.
  • $Y=\overline{B\setminus\iota([0,1]\times D^{2})}$.
  • $Y$ is not a solid torus.
Note that the complement of a knotted hole ball in $X$ is the connected sum of $X$ and $S^{1}\times D^{2}$.
@@
Since $\pi_{1}(\mathbb{R}\mathbb{P}^{3})\cong\mathbb{Z}/2\mathbb{Z}$ is finite, any torus $T$ in $\mathbb{R}\mathbb{P}^{3}$ is compressible. A sphere $S$ is obtained by compressing $T$ along a compressing disk $D$. The sphere $S$ bounds a ball $B$ since $\mathbb{R}\mathbb{P}^{3}$ is irreducible. If $B$ and $D$ are disjoint, then $T$ bounds a solid torus. Otherwise, $T$ bounds the complement of a 1-handle in $B$, which is a solid torus or a knotted hole ball. ∎
Let $L$ be a non-split link in $\mathbb{R}\mathbb{P}^{3}$. We say that a JSJ piece $M$ of the complement of $L$ is outermost if $\overline{\mathbb{R}\mathbb{P}^{3}\setminus M}$ is a disjoint union of solid tori and knotted hole balls. This definition can be applied to a non-split link in $S^{3}$ or $S^{1}\times D^{2}$. However, every JSJ piece for a non-split link in $S^{3}$ is outermost. The outermost piece for a non-split link in $S^{1}\times D^{2}$ contains the boundary of $S^{1}\times D^{2}$, and so it is unique. In any case, an outermost piece can be re-embedded as a link complement by reattaching the 1-handles adjacent to the knotted hole balls as in [9, Proposition 2.2] and [24, Lemma 3.6].
@@
  • If $T$ bounds two solid tori (i.e., $T$ is a Heegaard torus), then $e$ is unoriented.
Note that $T$ cannot bound a knotted hole ball to one side and a solid torus to the other side, for otherwise the ambient space would be $S^{3}$. Hence the orientation is well defined.
@@
Since $G_{L}$ is a tree, such a function $f$ exists. Suppose that a vertex $v\in V_{L}$ attains a local minimum of $f$. Then each edge incident to $v$ is oriented from $v$ or unoriented. Let $M$ denote the JSJ piece corresponding to $v$. Each boundary component of $M$ is a JSJ torus or the boundary of a neighborhood of a component of $L$. The orientations of edges imply that each boundary component of $M$ bounds a solid torus or a knotted hole ball disjoint from $M$. Hence $M$ is outermost. Since $G_{L}$ is finite, the minimum of $f$ exists. ∎
For a non-split link $L\subset\mathbb{R}\mathbb{P}^{3}$, let $\widetilde{L}\subset S^{3}$ denote the preimage of $L$ by the covering map $\pi\colon S^{3}\to\mathbb{R}\mathbb{P}^{3}$. By [24, Lemma 2.11] and the fact that there exist no embedded Klein bottles in $\mathbb{R}\mathbb{P}^{3}$ [8, 17], the JSJ tori of the complement of $\widetilde{L}$ is the preimage of the JSJ tori of the complement of $L$ by the covering map $\pi$. An outermost JSJ piece for $L$ is characterized as follows.
@@
Since a knotted hole ball is contained in a 3-ball $B$, the preimage of a knotted hole ball in $\mathbb{R}\mathbb{P}^{3}$ consists of its two copies, which are contained in the preimage of $B$. The preimage of a solid torus in $\mathbb{R}\mathbb{P}^{3}$ is one or two solid tori.
Suppose that $M$ is outermost. In other words, each component of $\overline{\mathbb{R}\mathbb{P}^{3}\setminus M}$ is a solid torus or a knotted hole ball. Assume that $\widetilde{M}$ is disconnected. Then the preimage of each component of the boundary $\partial\widetilde{M}$ consists of its two copies. Hence the preimage of each solid torus component of $\overline{\mathbb{R}\mathbb{P}^{3}\setminus M}$ also consists of its two copies. Since the preimage of each knotted hole ball must also consist of its two copies, the union $S^{3}$ of $\widetilde{M}$, the solid tori, and the knotted hole balls would be disconnected. Hence $\widetilde{M}$ is connected. Since the components of $\overline{S^{3}\setminus\widetilde{M}}$ that are not solid tori are the preimage of knotted hole ball components of $\overline{\mathbb{R}\mathbb{P}^{3}\setminus M}$, their number is even.
Conversely, suppose that $M$ is not outermost. Then Lemma 4.2 implies that $M$ is contained in a solid torus or a knotted hole ball. If $M$ is contained in a knotted hole ball, the preimage $\widetilde{M}$ is disconnected. Suppose that there is a JSJ torus $T\subset\partial M$ such that $M$ is contained in a solid torus $S$ bounded by $T$. Since an embedded torus in a solid torus bounds a solid torus or a knotted hole ball [24, Lemma 3.3], the components of $\overline{S\setminus M}$ consist of solid tori and knotted hole balls. If $T$ is a Heegaard torus of $\mathbb{R}\mathbb{P}^{3}$, then $M$ is outermost. If $\widetilde{M}$ is connected, and $T$ is not a Heegaard torus, then the components of $\overline{S^{3}\setminus\widetilde{M}}$ that are not solid tori consist of
@@
and so their number is odd. Note that if an orientable 3-manifold $N$ is not a solid torus, a finite cover of $N$ cannot be a solid torus. ∎
@@
The argument for split links is identical to that in the proof of [24, Theorem 4.5]. Suppose that a link $L_{0}$ maximally splits to $L_{0}^{\prime},L_{0,1},\dots,L_{0,m}$, where $L_{0}^{\prime}$ is a (possibly empty) non-split link. Then $\widetilde{L_{0}}=\pi^{-1}(L_{0})$ maximally splits to $\widetilde{L_{0}^{\prime}}=\pi^{-1}(L_{0}^{\prime})$ and $L_{0,j}^{k}$ for $1\leq j\leq m$ and $k=1,2$, where $L_{0,j}^{k}$ is a copy of $L_{0,j}$ as a link in $S^{3}$. The same applies to $L_{1}$. Since $\widetilde{L_{0}}$ and $\widetilde{L_{1}}$ are isotopic, their split summands are isotopic by Lemma 4.1. Hence we may assume that $L_{0}$ and $L_{1}$ are non-split links.
We consider the JSJ decompositions of the complements of $L_{0}$, $L_{1}$, and $\widetilde{L_{0}}\cong\widetilde{L_{1}}$. As mentioned above, the JSJ decomposition of the complement of $\widetilde{L_{0}}\cong\widetilde{L_{1}}$ is the lifts of those of $L_{0}$ and $L_{1}$. By Lemma 4.3, there is an outermost JSJ piece $M_{0}$ of the complement of $L_{0}$. By Lemma 4.4, $\widetilde{M_{0}}=\pi^{-1}(M_{0})$ is a single JSJ piece of the complement of $\widetilde{L_{0}}$. There is a JSJ piece $M_{1}$ of the complement of $L_{1}$ such that $\widetilde{M_{1}}=\pi^{-1}(M_{1})$ is isotopic to $\widetilde{M_{0}}$. By Lemma 4.4, the piece $M_{1}$ is also outermost. Since $M_{0}$ and $M_{1}$ are outermost, they can be re-embedded as the complements of links in $\mathbb{R}\mathbb{P}^{3}$. Since the preimage of a knotted hole ball in $\mathbb{R}\mathbb{P}^{3}$ consists of its two copies in $S^{3}$, the pieces $\widetilde{M_{0}}$ and $\widetilde{M_{1}}$ are equivariantly re-embedded as link complements. Since $M_{0}$ and $M_{1}$ are hyperbolic or Seifert fibered, they are isotopic in $\mathbb{R}\mathbb{P}^{3}$ by Theorems 2.5 and 3.6.
We show that the isotopy between $M_{0}$ and $M_{1}$ extends to an isotopy between the complements of $L_{0}$ and $L_{1}$ using an isotopy between the complements of $\widetilde{L_{0}}$ and $\widetilde{L_{1}}$. Since $M_{0}$ and $M_{1}$ are outermost, the remaining regions consist of solid tori and knotted hole balls. Since the preimage of a knotted hole ball in $\mathbb{R}\mathbb{P}^{3}$ consists of its two copies in $S^{3}$, the isotopy extends to the knotted hole balls. By [24, Theorem 4.5] for links in the solid torus, the isotopy in $S^{3}$ induces an isotopy between links in the remaining solid tori contained in $\mathbb{R}\mathbb{P}^{3}$. Therefore $L_{0}$ and $L_{1}$ are isotopic. ∎
@@ Acknowledgements
The author is grateful to Yuya Koda, Yuka Kotorii, Sonia Mahmoudi, Elisabetta A. Matsumoto, and Yuta Nozaki for their helpful discussions. This work was supported by the World Premier International Research Center Initiative Program, International Institute for Sustainability with Knotted Chiral Meta Matter (WPI-SKCM2), MEXT, Japan, and JSPS Program for Forming Japan’s Peak Research Universities (J-PEAKS) Grant Number JPJS00420230011.